‹ Back to the paper
Solve the following.
Consider the two events A and B. The probability that at least one of A or B occurs is 0· 6. If the…
Consider the two events $A$ and $B$. The probability that at least one of $A$ or $B$ occurs is $0\cdot 6$. If the probability of both $A$ and $B$ occurring is $0\cdot 3$, then calculate $P(A') + P(B')$.
Answer
Answer
$P(A\cup B)=P(A)+P(B)-P(A\cap B)$, so $P(A)+P(B)=0.6+0.3=0.9$
$P(A')+P(B')=2-[P(A)+P(B)]=2-0.9$
Final answer: 1.1